OpenStudy (anonymous):

Integrate e^((-3x+2sinx)/6)

OpenStudy (anonymous):

Thanks. It could be. Any indications would be helpful for me to investigate.

OpenStudy (anonymous):

Wait. That 1/6 is part of the exponent?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

You could have just left it and put 1/e^6 outside the integral.

OpenStudy (anonymous):

exp ((-3x +2sinx)\div6)

OpenStudy (anonymous):

Oh yah.

OpenStudy (anonymous):

e^1/6 is the integral?

OpenStudy (anonymous):

No. That's just a constant you can pull out of the integral

OpenStudy (anonymous):

That still looks like a tough one.

OpenStudy (anonymous):

yep haha

OpenStudy (anonymous):

\[-1/2\exp ^{2sinx-3x}\]

OpenStudy (anonymous):

That where i get to but can't find out what to do with 2sinx

OpenStudy (anonymous):

Wolfram Alpha - http://www.wolframalpha.com/input/?i=integrate+e%5E%282*sinx+-+3*x%29

OpenStudy (anonymous):

daaamn. if wolfram cant do it then i'm not going to try

OpenStudy (anonymous):

Are you supposed to get the indefinite integral or maybe an approximation of a definite integral?

OpenStudy (anonymous):

I think that is simpler than it seems but I have missed a trick?

OpenStudy (anonymous):

I need to get an indefinite integral

OpenStudy (anonymous):

Ok. I may have found the issue. Does it make a difference if I have (2cosx-3)

OpenStudy (anonymous):

with it

OpenStudy (anonymous):

This function as written does not have a closed form answer. Where did you get this integral?

OpenStudy (anonymous):

this looks the same: http://openstudy.com/study#/updates/4faf997ee4b059b524fa839e

OpenStudy (aravindg):

so was i late?

OpenStudy (anonymous):

I need to find the general solution to the following in implicit form \[dy/dx = (2cosx -3) \exp ^{(1/6)(-3x+2sinx))} y ^{5/6} (y>0)\]

OpenStudy (anonymous):

aha! that is easier i think, we have a substitution.

OpenStudy (anonymous):

Of course that is easier. Why did you hide ( 2 cos x -3?

OpenStudy (anonymous):

\[\int\limits{y^{\frac{-5}{6}}}dy = \int\limits{(2cosx - 3)e^{\frac{1}{6} (-3x + 2sinx)}}dx\] u = -3x + 2sinx

OpenStudy (anonymous):

And I use the substitution method?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

@WonHungLo i think we were wrong when we said we could take that factor out earlier.

OpenStudy (anonymous):

Thanks everybody. I can now investigate all this and see how we get to where we got to. I will need time but I am learning! YES!

OpenStudy (anonymous):

@hmidante, please write down the whole question and not one part of it.